# Social contagion under hybrid interactions

### v1

@@ Abstract
Threshold-driven models and game theory are two fundamental paradigms for describing human interactions in social systems. However, in mimicking social contagion processes, models that simultaneously incorporate these two mechanisms have been largely overlooked. Here, we study a general model that integrates hybrid interaction forms by assuming that a part of nodes in a network are driven by the threshold mechanism, while the remaining nodes exhibit imitation behavior governed by their rationality (under the game-theoretic framework). Our results reveal that the spreading dynamics are determined by the payoff of adoption. For positive payoffs, increasing the density of highly rational nodes can promote the adoption process, accompanied by a hybrid phase transition. The degree of rationality can regulate the spreading speed, with less rational imitators slowing down the spread. We further find that the results are opposite for negative payoffs of adoption. This model may provide valuable insights into understanding the complex dynamics of social contagion phenomena in real-world social networks.
@@ III.1 Case of positive payoffs
Figure 2: (a) Final adoption density $\rho_{\infty}$ as a function of the fraction of imitators $r$ for different values of $\phi$. The dashed lines correspond to the theoretical results. (b) and (c) Time evolution of the fraction of the two types of nodes for $r=0.26$ and $0.28$, respectively, under $\phi=0.5$. The positions of the two selected values of $r$ are marked as stars on the blue line in (a). The other parameters in the simulations are chosen as $N=5000$, $z=12$, $b=1$, $c=0$, and $\beta=1000$. Results are averaged over $10^{4}$ realizations.
@@
However, for large values of $\phi$, there appear two tipping points. In such cases, the type-0 nodes need adequate adopter neighbors to change state, potentially requiring more imitators to provide “external forces” (i.e., increasing $m/k$) for them to cross the barrier. As the number of imitators in a network progressively increases, there will be a critical point $r_{c}^{1}$ above which a giant connected cluster of imitators emerges (a traditional percolation process). For random ER networks, $r_{c}^{1}=1/z$. This giant cluster is vulnerable to external disturbances: once one node in this cluster becomes adopter, it will induce all the remaining nodes in it to change state through the imitation process, whose transient time is determined by the “rationality” of nodes $\beta$ [see Eq. (3)]. It should be emphasized that this cluster is different from the “vulnerable” cluster defined by Watts in Watts:2002 , which is composed of nodes with degree $k\leq 1/\phi$, and its transition from susceptible to adopter state (all nodes in the cluster become adopter) is almost instantaneous. Further increasing $r$, the size of the giant cluster of imitators will continuously increase, contributing to the linear growth of the final adoption density $\rho_{\infty}$. In this stage (before arriving the second tipping point), the imitation process dominates, as indicated by few type-0 nodes transitioning from susceptible to adopter state (see Fig. 2b).
@@ IV Theoretical Analysis
V Conclusion
Motivated by the observation that agents in real social networks exhibit diverse characteristics, we have proposed a simple networked contagion model that incorporates two kinds of nodes, each following its own state transition rule. Specifically, we assume that some nodes are sensitive to peer pressure and switch their state according to the threshold condition, while the remaining nodes behave as imitators, following the game-based formalism. We have introduced a minimalist game framework, assuming that (i) all adopters have a uniform payoff, and (ii) susceptible imitators follow the strategy of one of their neighbors at random according to the Fermi function. Our results show that the spreading dynamics depend strongly on the adoption payoffs. For positive payoffs, rational imitators are more likely to follow the decisions of the adopters. In this scenario, increasing the fraction of imitators in the system can effectively facilitate the cascading process, potentially leading to hybrid phase transitions. The speed of spreading is regulated by the rationality of the imitators: the more rational the nodes, the faster the growth of adopters. However, for negative payoffs, the result is the opposite. Highly rational nodes will play an inhibitory role in the cascade dynamics. Furthermore, as rationality decreases, the adoption density speeds up. Our model extends the current theoretical framework in social contagion modeling, which may provide deep insights for understanding various complex social contagion phenomena in reality.

### v2

@@ Abstract
Threshold-driven models and game theory are two fundamental paradigms for describing human interactions in social systems. However, in mimicking social contagion processes, models that simultaneously incorporate these two mechanisms have been largely overlooked. Here, we study a general model that integrates hybrid interaction forms by assuming that a part of nodes in a network are driven by the threshold mechanism, while the remaining nodes exhibit imitation behavior governed by their rationality (under the game-theoretic framework). Our results reveal that the spreading dynamics are determined by the payoff of adoption. For positive payoffs, increasing the density of highly rational nodes can promote the adoption process, accompanied by a double phase transition. The degree of rationality can regulate the spreading speed, with less rational imitators slowing down the spread. We further find that the results are opposite for negative payoffs of adoption. This model may provide valuable insights into understanding the complex dynamics of social contagion phenomena in real-world social networks.
@@ III.1 Case of positive payoffs
Figure 2: (a) Final adoption density $\rho_{\infty}$ as a function of the fraction of imitators $r$ for different values of $\phi$. The dashed lines correspond to the theoretical results. (b) Susceptibility $\chi$ as a function of $r$ for $\phi=0.6$. (c) and (d) Time evolution of the fraction of the two types of nodes for $r=0.28$ and $0.26$, respectively, under $\phi=0.5$. The positions of the two selected values of $r$ are marked as stars on the blue line in (a). The other parameters in the simulations are chosen as $N=5000$, $z=12$, $b=1$, $c=0$, and $\beta=1000$. Results are averaged over $10^{4}$ realizations.
@@
However, for large values of $\phi$, there appear two tipping points, suggesting a double phase transition Min:2018 ; Simon:2014 . To confirm this, we introduce a modified susceptibility defined as Souza:2019 ; Ferreira:2012 ; Arruda:2023
$\chi=N\frac{\langle\rho_{\infty}^{2}\rangle-\langle\rho_{\infty}\rangle^{2}}{\langle\rho_{\infty}\rangle}.$ (5)
A peak in the susceptibility $\chi$ signals the presence of a phase transition, and its position provides an estimate of the threshold. Taking $\phi=0.6$ as an example, we observe two distinct peaks in the $\chi(r)$ curve, with these peaks becoming more pronounced as the network size $N$ increases (see Fig. 2b). These findings confirm that the double phase transition is genuine.
To understand this phenomenon, notice that for large $\phi$, the type-0 nodes need a sufficient number of adopter neighbors to change state, potentially requiring more imitators to provide “external forces” (i.e., increasing $m/k$) for them to cross the barrier. As the number of imitators in a network progressively increases, there will be a critical point $r_{c}^{1}$ above which a giant connected cluster of imitators emerges (a traditional percolation process). For random ER networks, $r_{c}^{1}=1/z$. This giant cluster is vulnerable to external disturbances: once one node in this cluster becomes adopter, it will induce all the remaining nodes in it to change state through the imitation process, whose transient time is determined by the “rationality” of nodes $\beta$ [see Eq. (3)]. It should be emphasized that this cluster is different from the “vulnerable” cluster defined by Watts in Watts:2002 , which is composed of nodes with degree $k\leq 1/\phi$, and its transition from susceptible to adopter state (all nodes in the cluster become adopter) is almost instantaneous. Further increasing $r$, the size of the giant cluster of imitators will continuously increase, contributing to the linear growth of the final adoption density $\rho_{\infty}$. In this stage (before arriving the second tipping point), the imitation process dominates, as indicated by few type-0 nodes transitioning from susceptible to adopter state (see Fig. 2d).
@@ IV Theoretical Analysis
V Discussion
Motivated by the observation that agents in real social networks exhibit diverse characteristics, we have proposed a simple networked contagion model that incorporates two kinds of nodes, each following its own state transition rule. Specifically, we assume that some nodes are sensitive to peer pressure and switch their state according to the threshold condition, while the remaining nodes behave as imitators, following the game-based formalism. We have introduced a minimalist game framework, assuming that (i) all adopters have a uniform payoff, and (ii) susceptible imitators follow the strategy of one of their neighbors at random according to the Fermi function. Our results show that the spreading dynamics depend strongly on the adoption payoffs. For positive payoffs, rational imitators are more likely to follow the decisions of the adopters. In this scenario, increasing the fraction of imitators in the system can effectively facilitate the cascading process, potentially leading to double phase transitions. The speed of spreading is regulated by the rationality of the imitators: the more rational the nodes, the faster the growth of adopters. However, for negative payoffs, the result is the opposite. Highly rational nodes will play an inhibitory role in the cascade dynamics. Furthermore, as rationality decreases, the adoption density speeds up.
It is important to emphasize that incorporating the game-based formalism into the complex social contagion process offers several distinct advantages. First, it effectively captures realistic factors like individual rationality and adoption payoffs, which are typically crucial in real-world decision-making processes. Second, the model is highly flexible, accommodating different types of individuals—such as blocked nodes—simply by adjusting its parameters. Finally, it enables control over the evolution of the process without affecting the system’s asymptotic state, aligning well with observational data Iniguez:2018 ; Karsai:2016 . Our model extends the existing theoretical framework for social contagion modeling, which may offer valuable insights into understanding complex social contagion phenomena in reality.